3. a. Determine the equivalent resistance of the circuit b. Determine the current through points a, b, c, d and e. (L, I. eld, and le) c. How long does this whole circuit take to completely drain a AAA battery that has 6,210 Joule energy stored in it?

Please anawer all the parts of the question
Answer #1

To calculate equivalent resistance i just have used simple formula for series and parallel combinations. For series R = R1 + R2

for parallel 1/R = 1/R1 + 1/R2

20 Ohm and 10 Ohm in series and then this combination is parallel to 10 Ohm between d and e

To find all currents Ohm's law can be used.

In parallel combination the potentials across ends is same therefore potential across points b and c is also 1.5 V and potential across points d and e is also 1.5 V

Battery will be completely drained when it losses all energy in the dissipation into heat. Therefore i have used formula for heat .

Solution and calculation is attached on the scanned sheet.

From b oc cenneced in series therefore sesishance between b and c are 202 on Total Gaent2 Ampere ence- la = 012. Ampere Ampe Note Potential b ere OJS Ampe re 1 points ba bebween d ande Answer 3,6, 20,2 A Let cinuit take t seconds ·to、ambletely drain the AAA batfery means H t = a0700 se Answer 3., ciruit takes 5.7s hsurs to dvoin the battent ㎝blefel